# for Large Language Models

### v1

@@ Abstract
Hanxing Ding1,2thanks:   Equal contributions. Shuchang Tao311footnotemark: 1  Liang Pang1thanks:   Corresponding author Zihao Wei1,2 Jinyang Gao3Bolin Ding3Huawei Shen1,2 Xueqi Cheng1,2 1Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences 2 University of Chinese Academy of Sciences  3 Alibaba Group {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
@@ 5 Conclusion
Ethics Statements
Our proposed method aims to improve tool learning and planning in structured environments by leveraging clear API documentation and a program-empowered planning framework. While the approach holds significant potential for enhancing task automation and efficiency, we recognize the ethical considerations associated with its application. First, the method is inherently dependent on the quality of input data, such as API documentation, which may contain biases or inaccuracies. Inaccurate documentation could propagate errors or lead to unintended outcomes, highlighting the need for careful evaluation and validation of input data. Second, the automation enabled by our approach may impact labor dynamics in specific domains by reducing reliance on human expertise for repetitive or highly structured tasks. We emphasize that the system is designed to assist rather than replace human operators, aiming to improve productivity and free up human capacity for more creative or complex decision-making tasks.

### v2

@@ Abstract
Hanxing Ding1,2thanks:   Equal contributions. Shuchang Tao311footnotemark: 1  Liang Pang1thanks:   Corresponding author Zihao Wei1,2 Jinyang Gao3Bolin Ding3Huawei Shen1,2 Xueqi Cheng1,2 1State Key Laboratory of AI Safety, Institute of Computing Technology, CAS 2 University of Chinese Academy of Sciences  3 Alibaba Group {dinghanxing18s, pangliang, weizihao22z, shenhuawei, cxq}@ict.ac.cn
@@ 5 Conclusion
Acknowledgement
This work was supported by the Strategic Priority Research Program of the CAS under Grants No.XDB0680302, the National Natural Science Foundation of China (NSFC) under Grants No62276248, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.